Find the equation to the straight line which passes through the point 5,6 and has intercept on the axis equal in magnitude but opposite in sign

Grade:10

1 Answers

Arun
4 years ago
Dear student 
 
if intercept are equal and opposite in sign then its slope whould be – 1
 
hence equation is x/ a – y/a = 1
 
5/a – 6/a = 1
 
hence a = – 1
 
hence equation is-
 
- x + y = 1